If 50 of them are taken as samples, what is the probability that their mean height is less than 150cm?



1
Expert's answer
2022-05-26T17:12:54-0400

Given "\mu=151cm, \sigma=6cm, n=50."

"P(X<150)=P(Z<\dfrac{150-151}{6/\sqrt{50}})"

"\approx P(Z<-1.1785)\approx 0.1193"